# Lossy Beyond Diagonal Reconfigurable Intelligent Surfaces: Modeling and Optimization

### v1

@@ V-A BD-RIS-aided SISO System
Figure 4: Average rate versus resistance $R$ for lossy BD-RIS with group- and forest-connected architectures. In the legend, ”MA” denotes MM-ADMM, ”LC” denotes low-complexity, ”Tri” stands for tridiagonal, ”Arr” stands for arrowhead, and ”UB” stands for upper-bound ($M=30$, $\bar{M}\in\{1,6,30\}$, $P=20$ dBm).
We first evaluate the performance of the proposed MM-ADMM algorithm and the low-complexity algorithm for BD-RIS-aided SISO systems by plotting the average rate versus resistance $R$ in Fig. 4. The rate for each channel realization is computed as $\log_{2}(1+P\frac{|h|^{2}}{\sigma^{2}})$, and the upper-bound for the maximum received signal power of group- and forest-connected BD-RISs is given by $\left(\sum_{g=1}^{G}\|\mathbf{h}_{RI,g}\|\|\mathbf{h}_{IT,g}\|+|h_{RT}|\right)^{2}$. From Fig. 4, we have the following observations.
First, for both algorithms, BD-RIS with a group-connected architecture outperforms D-RIS across the entire considered range of $R$. With a relatively small group size, e.g., $\bar{M}=6$, group-connected BD-RIS maintains superior performance over D-RIS, exhibiting only marginally faster degradation while delivering substantially better performance. This is because when the circuit complexity is relatively low, the enhanced wave manipulation flexibility provided by more reconfigurable admittance components of BD-RIS outweighs their losses. However, this does not hold for fully-connected cases since the quadratically increased admittances introduce significant losses. As a result, the fully-connected BD-RIS with $\bar{M}=M$ exhibits significantly faster performance degradation compared to D-RIS. This indicates that the losses in BD-RIS have the most significant impact on the fully-connected architecture, implying that the group-connected architecture with a proper group size is more practical. Although MM-ADMM algorithm outperforms the low-complexity algorithm for D-RIS, their performance converges for fully-connected BD-RIS when $R\leq 2\ \Omega$. Beyond this point, the low-complexity solution degrades faster, highlighting its approximation becomes less valid for both D-RIS and high-loss fully-connected BD-RIS.
Second, for both algorithms, forest-connected BD-RIS with both tridiagonal and arrowhead forms outperforms D-RIS across the entire considered range of $R$. Both algorithms show no significant performance difference between forest-connected BD-RIS with $\bar{M}=6$ and tree-connected BD-RIS with $\bar{M}=M$. This is because the tree-connected architecture introduces only a small number of additional tunable admittances compared to the forest-connected one. For the low-complexity algorithm, these extra tunable admittances help to compensate for approximation errors. As a result, the tridiagonal form of the tree-connected BD-RIS outperforms the forest-connected BD-RISs under the low-complexity algorithm. Notably, the performance gap between the low-complexity and MM-ADMM algorithm remains relatively small for all forest- and tree-connected architectures, which suggests that the low-complexity algorithm is more practical as it requires significantly lower computational complexity.
Third, when accounting for lossy tunable admittance components, both algorithms demonstrate that group-connected BD-RIS (excluding the fully-connected architecture) outperforms forest-connected BD-RIS. This observation is not consistent with the conclusion for lossless cases, where the forest- and group-connected BD-RISs could achieve the same performance [10]. This is because, lossless admittances have purely imaginary and arbitrarily tunable admittance values, where $(2\bar{M}-1)G$ components are sufficient to reach the performance upperbound. However, in lossy cases, both architectures cannot reach the performance upperbound, while the former has more components to be tuned to compensate for the power loss.
@@
Figure 5: Average rate versus group size $\bar{M}$ for lossy BD-RIS with group- and forest-connected architectures. In the legend, ”MA” denotes MM-ADMM, ”LC” denotes low-complexity, ”Tri” stands for tridiagonal, ”Arr” stands for arrowhead, and ”UB” stands for upper-bound ($M=60$, $P=20$ dBm, $R=2.5\ \Omega$).
@@
Fig. 5 compares the performance of group-connected and forest-connected BD-RISs as the group size $\bar{M}$ increases, evaluating the MM-ADMM and low-complexity algorithms. The results reveal two interesting trends: First, for group-connected BD-RIS, the performance gap between two algorithms narrows as $\bar{M}$ increases, demonstrating that a larger group size provides greater flexibility to compensate for approximation errors. Second, for both algorithms, forest-connected BD-RIS exhibits only marginal performance improvements as $\bar{M}$ increases from 3 to 6. This occurs because the increase in $\bar{M}$ within $\{3,\ldots,6\}$ introduces only a small number of additional tunable admittances. In summary, a relative small group size (e.g., $\bar{M}\in\{4,5,6\}$) in the group-connected architecture with the low-complexity algorithm can achieve good performance while keeping computational complexity low.
V-B BD-RIS-aided MU-MISO System
@@
We then evaluate the sum-rate performance of the BD-RIS-aided MU-MISO system by plotting the sum-rate versus resistance $R$ in Fig. 6. The simulations employ the tridiagonal form of the forest-connected architecture, with both the number of users, $K$, and the number of antennas at the BS, $N$, fixed to 4. From Fig. 6, we have the following observations.
First, all BD-RIS architectures outperform D-RIS across the entire considered range of $R$. Unlike the SISO case where fully-connected BD-RIS achieves comparable performance to D-RIS at $R=3\ \Omega$, in MU-MISO case, fully-connected BD-RIS maintains significant gains over D-RIS and other architectures even under a large $R$. This enhanced performance likely stems from the proper precoder design effectively compensating for the huge losses at BD-RIS.
@@
In Fig. 7, we compare the sum-rate performance of group- and forest-connected BD-RISs versus the transmit power $P$. As expected, fully-connected BD-RIS achieves the best performance, and the group-connected BD-RIS outperforms forest-connected BD-RIS with the same group size. In contrast to single-user SISO systems, in MU-MISO systems, the presence of more interconnections in lossy BD-RISs is beneficial. This is because the multiuser MISO channel provides a higher degree of freedom (DoF) and relies more heavily on interconnections to exploit this increased DoF [12]. As a result, additional tunable admittances enhance performance, even in the presence of losses.
Figure 7: Sum-rate versus transmit power $P$ for lossy BD-RIS with group- and forest-connected architectures ($N=K=4$, $M=32$, $\bar{M}\in\{1,4,32\}$, $R=2.5\ \Omega$).
@@ VI Conclusion
In this paper, we investigate lossy BD-RIS modeling and optimization for SISO and MU-MISO systems. Specifically, we model the lossy reconfigurable admittance network by representing each tunable admittance component with a practical varactor based on a lumped circuit model. Taking into account the practical capacitance range of the varactor, each tunable admittance is constrained within a feasible range.
Building on the proposed lossy BD-RIS model, we optimize the scattering matrix for BD-RIS-aided SISO systems to maximize the received signal power. To solve this problem, we employ an MM framework to transform it into a more tractable form and then develop an ADMM algorithm to efficiently handle the complex constraints. To reduce computational demands, we further propose a low-complexity algorithm that approximates the original problem. Next, we extend the optimization to jointly design the transmit precoder and scattering matrix to maximize the sum-rate for BD-RIS-aided MU-MISO systems. This problem is reformulated as a multi-block optimization using FP theory, and an ADMM algorithm, similar to the SISO case, is applied to efficiently update the scattering matrix.
Under the proposed lossy BD-RIS model, simulations results show that all BD-RIS architectures outperform D-RIS in both SISO and MU-MISO systems. However, the optimal BD-RIS architectures in the lossless case are not necessarily optimal when losses are considered, e.g. fully-connected and tree-connected BD-RISs fail to achieve the best performance in SISO systems. This suggests that the characterization of the fundamental Pareto frontier between performance and complexity of lossless architectures in [26, 12] needs to be revisited in the lossy case.

### v2

@@ V-B BD-RIS-aided SISO System
Figure 7: Average rate versus resistance $R$ for lossy BD-RIS with group-connected architecture. In the legend, ”MA” denotes $M$-ADMM, ”LC” denotes low-complexity, and ”UB” stands for upper-bound ($M=30$, $\bar{M}\in\{1,6,30\}$, $P=20$ dBm).
(a) $M$-ADMM algorithm
(b) Low-complexity algorithm
@@
Figure 8: Average rate versus resistance $R$ for lossy BD-RIS with forest-connected architectures. In the legend, ”MA” denotes $M$-ADMM, ”LC” denotes low-complexity, ”Tri” stands for tridiagonal, ”Arr” stands for arrowhead, and ”UB” stands for upper-bound ($M=30$, $\bar{M}\in\{1,6,30\}$, $P=20$ dBm).
We first evaluate the performance of the proposed $M$-ADMM algorithm and the low-complexity algorithm for BD-RIS-aided SISO systems by plotting the average rate versus resistance $R$ in Figs. 7 and 8. The rate for each channel realization is computed as $\log_{2}(1+P\frac{|h|^{2}}{\sigma^{2}})$, and the upperbound [^12] The gap between the upperbound and the rate at $R=0\ \Omega$ arises because, at $R=0\ \Omega$, each tunable admittance remains constrained to the imaginary axis with bounded capacitance values $C_{m_{g},n_{g}}$. for the maximum received signal power of group- and forest-connected BD-RISs is given by $\left(\sum_{g=1}^{G}\|\mathbf{h}_{RI,g}\|\|\mathbf{h}_{IT,g}\|+|h_{RT}|\right)^{2}$. We plot the performance of both algorithms for group-connected BD-RIS in Fig. 7, from which we have the following observations.
First, for both algorithms, BD-RIS with a group-connected architecture outperforms D-RIS across the entire considered range of $R$. With a relatively small group size, e.g., $\bar{M}=6$, group-connected BD-RIS maintains superior performance over D-RIS, exhibiting only marginally faster degradation while delivering substantially better performance. This is because when the circuit complexity is relatively low, the enhanced wave manipulation flexibility provided by more reconfigurable admittance components of BD-RIS outweighs their losses. However, this does not hold for fully-connected cases since the quadratically increased admittances introduce significant losses. As a result, the fully-connected BD-RIS with $\bar{M}=M$ exhibits significantly faster performance degradation compared to D-RIS. This indicates that the losses in BD-RIS have the most significant impact on the fully-connected architecture, implying that the group-connected architecture with a proper group size is more practical.
Second, although the $M$-ADMM algorithm generally outperforms the low-complexity algorithm for D-RIS, their performance is nearly identical for fully-connected BD-RIS when $R\leq 2\ \Omega$ and group-connected BD-RIS with $\bar{M}=6$. For D-RIS, the average rate achieved by MM–ADMM exceeds that of the low-complexity solution by 7.12%, 6.60%, 6.29%, and 6.07% for $R=\{0,1,2,3\}\ \Omega$, respectively. For fully-connected case when $R=3\ \Omega$, the performance gap is 3.83%. These observations indicate that the approximation in the low-complexity algorithm becomes less accurate for D-RIS and for high-loss fully-connected BD-RIS.
In Fig. 8, the performance of both algorithms for forest-connected BD-RIS is presented, from which the following observations can be made.
First, for both algorithms, forest-connected BD-RIS with both tridiagonal and arrowhead forms outperform D-RIS across the entire considered range of $R$. Both algorithms show no significant performance difference between forest-connected BD-RIS with $\bar{M}=6$ and tree-connected BD-RIS with $\bar{M}=M$. This is because the tree-connected architecture introduces only a small number of additional tunable admittances compared to the forest-connected one. For the low-complexity algorithm, these extra tunable admittances help to compensate for approximation errors. As a result, the tridiagonal form of the tree-connected BD-RIS outperforms the forest-connected BD-RIS under the low-complexity algorithm.
Second, the performance gap between the low-complexity and $M$-ADMM algorithm remains relatively small for all forest- and tree-connected architectures. For example, for the tridiagonal tree-connected BD-RIS, $M$-ADMM outperforms the low-complexity solution by only 1.71%, 1.33%, 1.31%, and 1.41% for $R=\{0,1,2,3\}\ \Omega$, respectively, indicating that the low-complexity algorithm is well suited to forest-connected architectures given its much lower computational complexity.
In both Figs. 7 and 8, when accounting for lossy tunable admittances, both algorithms demonstrate that group-connected BD-RIS (excluding the fully-connected architecture) outperforms forest-connected BD-RIS. This observation is not consistent with the conclusion for lossless cases, where the group- and forest-connected BD-RISs could achieve the same performance [nerinigraphtheory]. This is because, lossless admittances have purely imaginary and arbitrarily tunable admittance values, where $(2\bar{M}-1)G$ components are sufficient to reach the performance upperbound. However, in lossy cases, both architectures cannot reach the performance upperbound, while the former has more components to be tuned to compensate for the power loss.
Figure 9: Average rate versus group size $\bar{M}$ for lossy BD-RIS with group- and forest-connected architectures. In the legend, ”MA” denotes $M$-ADMM, ”LC” denotes low-complexity, ”Tri” stands for tridiagonal, ”Arr” stands for arrowhead, and ”UB” stands for upper-bound ($M=60$, $P=20$ dBm, $R=2.5\ \Omega$).
@@
Fig. 9 compares the performance of group-connected and forest-connected BD-RISs as the group size $\bar{M}$ increases, evaluating the $M$-ADMM and low-complexity algorithms. The results reveal two interesting trends: First, for group-connected BD-RIS, the performance gap between two algorithms narrows as $\bar{M}$ increases, demonstrating that a larger group size provides greater flexibility to compensate for approximation errors. Second, for both algorithms, forest-connected BD-RIS exhibits only marginal performance improvements as $\bar{M}$ increases from 3 to 6. This occurs because the increase in $\bar{M}$ within $\{3,\ldots,6\}$ introduces only a small number of additional tunable admittances. In summary, a relative small group size (e.g., $\bar{M}\in\{4,5,6\}$) in the group-connected architecture with the low-complexity algorithm can achieve good performance while keeping computational complexity low.
V-C BD-RIS-aided MU-MISO System
@@
We then evaluate the sum-rate performance of the BD-RIS-aided MU-MISO system by plotting the sum-rate versus resistance $R$ in Fig. 10. The simulations employ the tridiagonal form of the forest-connected architecture, with both the number of users, $K$, and the number of antennas at the BS, NN, fixed to 4. From Fig. 10, we have the following observations.
First, all BD-RIS architectures outperform D-RIS across the entire considered range of $R$. Unlike the SISO case where fully-connected BD-RIS achieves comparable performance to D-RIS at $R=3\ \Omega$, in MU-MISO case, fully-connected BD-RIS maintains significant gains over D-RIS and outperforms other architectures even under a large $R$. This enhanced performance likely stems from the proper precoder design effectively compensating for the huge losses at BD-RIS.
@@
In Fig. 11, we compare the sum-rate performance of group- and forest-connected BD-RISs versus the transmit power $P$. As expected, fully-connected BD-RIS achieves the best performance, and the group-connected BD-RIS outperforms forest-connected BD-RIS with the same group size. The presence of more interconnections in lossy BD-RIS is beneficial in MU-MISO systems compared to single-user SISO systems. This is because the multiuser MISO channel provides a higher degree of freedom (DoF) and relies more heavily on interconnections to exploit this increased DoF [band_stem_zheyu]. As a result, additional tunable admittances enhance performance, even in the presence of losses.
In Fig. 12, we plot the sum-rate performance versus circuit complexity for the group-connected BD-RIS. Unlike the Pareto frontier of the lossless BD-RIS reported in [band_stem_zheyu], increasing the circuit complexity does not always lead to higher performance; for example, the fully-connected architecture fails to achieve the best performance when $R\geq 2\ \Omega$. Under significant losses (i.e., under large $R$), adding more tunable components may in fact degrade the overall performance. This observation highlights the importance of carefully balancing performance and complexity, emphasizing the need to select an appropriate level of complexity for lossy BD-RIS.
Figure 11: Sum-rate versus transmit power $P$ for lossy BD-RIS with group- and forest-connected architectures. In the legend, ”Tri” stands for tridiagonal, and ”Arr” stands for arrowhead ($N=K=4$, $M=32$, $\bar{M}\in\{1,4,32\}$, $R=2.5\ \Omega$).
Figure 12: Performance-complexity trade-off for lossy BD-RIS with a group-connected architecture, given corresponding group size $\bar{M}$ and different values of $R$ ($N=K=4$, $M=32$, $P=20$ dBm).
@@ VI Conclusion
In this paper, we investigate lossy BD-RIS modeling and optimization for SISO and MU-MISO systems. Specifically, we model the lossy reconfigurable admittance network by representing each tunable admittance component with a practical varactor diode based on a lumped circuit model. Taking into account the practical capacitance range of the varactor, each tunable admittance is constrained within a feasible range.
Building on the proposed lossy BD-RIS model, we optimize the scattering matrix for BD-RIS-aided SISO systems to maximize the received signal power. To solve this problem, we employ an $M$ framework to transform it into a more tractable form and then develop an ADMM algorithm to efficiently handle the complex constraints. To reduce computational demands, we further propose a low-complexity algorithm that approximates the original problem. Next, we extend the optimization to the joint design of the transmit precoder and scattering matrix to maximize the sum-rate for BD-RIS-aided MU-MISO systems. This problem is reformulated as a multi-block optimization using FP theory, and an ADMM algorithm, similar to the SISO case, is applied to efficiently update the scattering matrix.
Under the proposed lossy BD-RIS model, simulations results show that all BD-RIS architectures outperform D-RIS in both SISO and MU-MISO systems. However, the optimal BD-RIS architectures in the lossless case are not necessarily optimal when losses are considered, e.g. fully-connected and tree-connected BD-RISs fail to achieve the best performance in SISO systems. This suggests that the characterization of the fundamental Pareto frontier between performance and complexity for lossless architectures in [parteo_frontier_nerini, band_stem_zheyu] needs to be revisited in the lossy case. A thorough investigation of the performance–complexity trade-off for lossy BD-RIS architectures is thus left as an insightful future work.
[Explicit forms of $\mathbf{P}$ and $U$] i) Group-connected architecture: $U=\frac{\bar{M}(\bar{M}+1)}{2}$ and $\bar{\mathbf{y}}_{g}$ collects the $\bar{M}$ diagonal and $\frac{\bar{M}(\bar{M}-1)}{2}$ upper triangular entries of $\bar{\mathbf{Y}}_{g}$. The mapping matrix $\mathbf{P}\in\{0,1\}^{\bar{M}^{2}\times\frac{\bar{M}(\bar{M}+1)}{2}}$ is given by
$[\mathbf{P}]_{\bar{M}(i-1)+j,k}=$ (62)
[…]
$\forall i,j\in\mathcal$$\bar${M}}.
ii) Forest-connected architecture (tridiagonal form): $U=2\bar{M}-1$ and $\bar{\mathbf{y}}_{g}$ collects the $\bar{M}$ diagonal entries and $\bar{M}-1$ off-diagonal entries of $\bar{\mathbf{Y}}_{g}$ based on (5). The mapping matrix $\mathbf{P}\in\{0,1\}^{\bar{M}^{2}\times(2\bar{M}-1)}$ is given by
$[\mathbf{P}]_{\bar{M}(i-1)+j,k}=$ (63)
[…]
$\forall i,j\in\mathcal$$\bar${M}}.
iii) Forest-connected architecture (arrowhead form): $U=2\bar{M}-1$ and $\bar{\mathbf{y}}_{g}$ collects the $\bar{M}$ diagonal entries and $\bar{M}-1$ off-diagonal entries of $\bar{\mathbf{Y}}_{g}$ based on (6). The mapping matrix $\mathbf{P}\in\{0,1\}^{\bar{M}^{2}\times(2\bar{M}-1)}$ is given by
[…]
$\forall i,j\in\mathcal$$\bar${M}}.